OpenStudy (anonymous):

HELP PLEASE :((( Input the slope and y -intercept without solving for y . Reduce all fractional answers to lowest terms. 2x + 3y = 6 m = b =

OpenStudy (calculusxy):

Y=mx+b Why don't you just replace the numbers for the positions in the formula. Then simplify it.

OpenStudy (anonymous):

3y=-2x+6 y=-2/3x+6/3 y=-2/3x+2 m=-2/3 b=2